# Exploring the generalizability of the optimal 0.234 acceptance rate in random-walk Metropolis and parallel tempering algorithms

### v1

@@ Abstract
For random-walk Metropolis (RWM) and parallel tempering (PT) algorithms, an asymptotic acceptance rate of around 0.234 is known to be optimal in the high-dimensional limit. Yet, the practical relevance of this value is uncertain due to the restrictive conditions underlying its derivation. We synthesise previous theoretical advances in extending the 0.234 acceptance rate to more general settings, and demonstrate the applicability and generalizability of the 0.234 theory for practitioners with a comprehensive empirical simulation study on a variety of examples examining how acceptance rates affect Expected Squared Jumping Distance (ESJD). Our experiments show the optimality of the 0.234 acceptance rate for RWM is surprisingly robust even in lower dimensions across various proposal and multimodal target distributions which may or may not have an i.i.d. product density. Experiments on parallel tempering also show that the idealized 0.234 spacing of inverse temperatures may be approximately optimal for low dimensions and non i.i.d. product target densities, and that constructing an inverse temperature ladder with spacings given by a swap acceptance of 0.234 is a viable strategy. However, we observe the applicability of the 0.234 acceptance rate heuristic diminishes for both RWM and PT algorithms below a certain dimension which differs based on the target density, and that inhomogeneously scaled components in the target density further reduces its applicability in lower dimensions.
@@ 1 Introduction
Markov chain Monte Carlo (MCMC) methods, such as the Metropolis algorithm, are used to draw samples from complex high-dimensional target probability distributions [2, 19]. They enjoy strong theoretical asymptotic guarantees of accuracy, converging to the target distribution in stationarity [2, 19]. However, they may also be inefficient and take an intractable amount of time to provide a good approximation of the target distribution they are sampling [22]. Therefore, the analysis of their running time is an important practical issue to consider.
@@
[20] proved an important result: in the state space high-dimensional limit $d\to\infty$, the asymptotic acceptance rate of a Metropolis algorithm is central to its efficiency, and that for a target distribution that has density $\pi$ with respect to Lebesgue measure on $\mathbb{R}^{d}$ where $\pi(\mathbf{x})=\prod_{i=1}^{d}f(x_{i})$ has an i.i.d. product form, and where $f$ is a sufficiently smooth single-dimensional density function such that $\log f$ is continuously differentiable and $\mathbb{E}_{f}\left[\left(\frac{f^{\prime}(X)}{f(X)}\right)^{2}\right]<\infty$, an asymptotic acceptance rate of approximately 0.234 maximises the efficiency of a random-walk Metropolis (RWM) algorithm with a Gaussian increment proposal distribution starting at stationarity. However, these assumptions are very restrictive. Many practical scenarios often involve finite and relatively small dimensions. Additionally, many probability distributions do not have densities of the i.i.d. product form, and for those that do, sampling a one-dimensional target is much simpler than using RWM. A burgeoning active research area is dedicated to showing that this figure still applies beyond these assumptions, and much of this focuses on the i.i.d. product assumption.
@@
Many papers have shown that the optimal average acceptance rate of 0.234 for RWM holds greater generality than the assumptions used in [20]. These extensions include: non-Gaussian proposals [17], discrete hypercube target distributions [21], Gibbs random fields [3], independent target components with inhomogeneous scaling [5, 4, 23], partial updates where not all components are updated at once [16], elliptical symmetric unimodal target densities [28], and infinite-dimensional target distributions with non-trivial dependence structures [14]. The 0.234 acceptance rate in optimal scaling has also been shown to apply to the spacing and swapping between inverse temperatures in the parallel tempering MCMC method [1, 26]. That being said, there are some cases in which the 0.234 case may not hold, such as in the case of discontinuous targets [18] where the optimal acceptance rate is approximately 0.1353. Additionally, these results of optimality are still theoretical ideals, with similar assumptions to the product i.i.d. form and in the limit of $d\to\infty$ where the Markov chain converges to a diffusion process.
@@
We begin our paper with a description of the RWM and its optimal scaling framework in Section 2. In Section 3, we describe RWM experiments in low dimensions that show where an acceptance rate of 0.234 can be optimal in the sense that it maximises the Expected Squared Jumping Distance (ESJD) for various proposal and target densities which are not necessarily all i.i.d. product forms. Next, we describe the parallel tempering method and its own optimal scaling framework in Section 4, and show using experiments in Section 5 that the 0.234 swap acceptance rate in parallel tempering may also be optimal in lower dimensions for the multivariate Gaussian target density and a Gaussian mixture target density which is not of an i.i.d. product form. Lastly, we discuss the implications of our findings in Section 6.
@@ 2 Random-walk Metropolis Background
We first introduce the Metropolis algorithm and the Optimal Scaling framework. We then discuss measures to evaluate the efficiency of MCMC algorithms, such as the expected squared jumping distance, and explain how the asymptotic acceptance rate of a MCMC algorithm is tied to the algorithm’s efficiency.
@@ 2.1 Metropolis Algorithm
A Metropolis algorithm [15] constructs a Markov chain $\mathbf{x}^{(0)},\mathbf{x}^{(1)},\dots$ that converges to the target probability distribution $\pi$ in the long-term after many runs, i.e. the Markov chain has a stationary distribution equal to the target distribution. When taking each step in the Markov chain, we generate a new state $\mathbf{y}=\mathbf{x}^{(t)}+\epsilon$ where $\epsilon$ is sampled from an increment “proposal” density $Q$ having $Q(\epsilon)=Q(-\epsilon)$ for all $\epsilon$ in its domain and accept this new state with probability $\alpha(\mathbf{x},\mathbf{y})$ given by $\alpha=\min\left\{1,\dfrac{\pi(\mathbf{y})}{\pi(\mathbf{x})}\right\}$. If the proposed value is accepted, we set the next state $\mathbf{x}^{(t+1)}=\mathbf{y}$. Otherwise, we set $\mathbf{x}^{(t+1)}=\mathbf{x}^{(t)}$.
@@ 2.2 Optimal Scaling Framework
Optimal Scaling [20, 23] is one of the most successful frameworks for performing asymptotic analysis of high-dimensional MCMC methods, and provides mathematically-grounded guidance on how to best optimise MCMC performance by tuning the “scaling” parameter(s) of the proposal distribution. Typically, a multivariate Gaussian distribution with mean 0 is used as the increment proposal distribution, and the proposal’s scaling is its variance $\sigma^{2}$. If the proposal’s variance is too small, too many proposed steps are accepted by the algorithm, but each step does not explore the state space much. Vice versa, if the proposal’s variance is too large, too few proposed steps are accepted, leading to slow exploration of the state space since the algorithm stays at a state for too long. Therefore, the proposed variance is crucial to the performance of the algorithm; the optimal scaling should balance the exploration of new areas in the state space with exploiting high-density areas of the target distribution. Different Metropolis algorithms have shown to have different optimal choices of the proposal distribution variance $\sigma^{2}$.
@@
We next discuss how to evaluate the efficiency of an MCMC algorithm for determining optimal scaling. A very popular measure of efficiency is the Expected Squared Jumping Distance (ESJD) metric [9, 30] which we use for our experiments. The expected squared jumping distance measures how far, in expectation, the MCMC chain moves in a single iteration. For the standard random-walk Metropolis algorithm, we define this as
$\mathbb{E}\left[\left\|x^{(t+1)}-x^{(t)}\right\|^{2}\right]\approx\frac{1}{n-1}\sum_{i=1}^{n-1}\left\|x^{(t+1)}-x^{(t)}\right\|^{2}$
@@
where $n$ is the total number of iterations of the algorithm. Maximising the ESJD aligns with minimizing the first-order auto-correlation of the Markov chain and subsequently maximises efficiency if the higher-order auto-correlations are monotonically increasing relative to the first-order auto-correlation [9, 30].
@@
There are other notions of efficiency of a Markov chain, but in the high-dimensional limit $d\to\infty$ where the chain converges to a diffusion process, all efficiency measures are effectively equivalent [20, 23]. A key result of these referenced papers is that, given a Metropolis algorithm with a Gaussian increment proposal distribution $Q=N(0,\frac{\ell^{2}}{d}I_{d})$ where $\ell>0$ is a fixed scaling constant and $I_{d}$ is the identity matrix, maximising the algorithm’s speed measure $h(\ell)$, which is a function of the scaling constant $\ell$, yields the most efficient asymptotic diffusion. Furthermore, the speed measure has a clear relation to a much simpler quantity to estimate: the asymptotic acceptance rate of the proposed new states (moves) of the algorithm, defined as
@@
Both the speed measure $h(\ell)$ and asymptotic acceptance rate $a(\ell)$ are functions of the scaling constant $\ell$. As mentioned in Section 1, when the target distribution that has density $\pi$ with respect to Lebesgue measure on $\mathbb{R}^{d}$ has an i.i.d. product form $\pi(\mathbf{x})=\prod_{i=1}^{d}f(x_{i})$, where $f$ is a sufficiently smooth single-dimensional density function such that $\log f$ is continuously differentiable and $\mathbb{E}_{f}\left[\left(\frac{f^{\prime}(X)}{f(X)}\right)^{2}\right]<\infty$, for a random-walk Metropolis algorithm with a Gaussian increment proposal distribution starting at stationarity, the scaling constant $\ell$ that maximises the speed measure $h(\ell)$ corresponds to an asymptotic acceptance rate $a(\ell)$ of approximately 0.234 [20, 23]. Hence, an asymptotic acceptance rate of approximately 0.234 should necessarily optimise a measure of efficiency such as the ESJD of the algorithm.
3 Random Walk Metropolis Simulations
@@
In this section, we explore the applicability of the 0.234 acceptance rate figure for RWM algorithms using several practical RWM examples in lower dimensions. We begin with simple examples where the target densities meet all the key assumptions required for the 0.234 acceptance rate to be optimal, as established in the literature. Next, we investigate examples that do not use Gaussian proposal distributions, and then we explore other densities that may lack key assumptions, such as smoothness or the i.i.d. product form. We estimate the Expected Squared Jumping Distance (ESJD) of the algorithm as a metric of the algorithm’s efficiency as mentioned in Section 2.2 and aim to see under what conditions an acceptance rate of approximately 0.234 maximises the ESJD.
Our experiments show which acceptance rate value (which is implicitly determined by the proposal variance) maximises the ESJD for these various target distributions in lower dimensions. It is important to note that functions of acceptance rates are somewhat flat around the optimal values and that the values and figures reported are still prone to random error inherent in a Monte Carlo simulation. To reduce the random error, we run each experiment over 5 different seeds and average the results. In each experiment, we run RWM simulations with 40 different variance values using a Gaussian proposal (unless otherwise mentioned) for 100,000 runs each and recorded the ESJD and acceptance rate estimates for the simulations. For each example, we conduct this experiment over multiple dimension sizes. Overall, each example has 200 simulations total over 5 seeds, and each simulation runs for 100,000 iterations.
@@
We begin with simpler examples demonstrating that the 0.234 acceptance rate is still roughly optimal for lower dimensions with any i.i.d. product target distribution $\pi(\mathbf{x})=\prod_{i=1}^{d}f(x_{i})$ and a Gaussian proposal. As mentioned in Section 2, [20] proved the 0.234 optimal acceptance rate for i.i.d. product target distributions with a standard Gaussian proposal distribution in the infinite-dimensional limit, and [23] and [6] extended this with MCMC simulations showing that with a i.i.d. product target distribution and Gaussian proposal, the 0.234 acceptance rate seems approximately optimal for dimensions as low as 10. Our experiment provides further evidence for this being the case, and we extend this with simulations for $d=2,5$.
We present two examples: the first example sets the singular dimensional component density $f$ to be Gamma$(3,2)$ with shape parameter 3 and scale parameter 2, and the second example uses Beta$(3,2)$ as the component density $f$ with shape parameters 3 and 2. We use a multivariate Gaussian with a mean of 0 as the increment proposal distribution in both cases. We run this experiment in dimensions 2, 5, 10, 30, and 50.
Figure 1: ESJD for various RWM acceptance rates for the i.i.d. gamma target distribution in dimensions $d=$ 2, 5, 10, 30, 50 (from top-left to bottom-right).
Figure 2: ESJD for various RWM acceptance rates for the i.i.d. beta target distribution in dimensions $d$ = 2, 5, 10, 30, 50 (from top-left to bottom-right).
Table 1: Experimental optimal RWM acceptance rates for i.i.d. gamma and beta target distributions in different dimensions. Numerical values are the observed experimental acceptance rates maximising the ESJD.
Dimension $d$
Target Density 2 5 10 30 50
i.i.d. $Gamma(3,2)$ 0.3009 0.2483 0.2095 0.2239 0.2338
i.i.d. $Beta(3,2)$ 0.3780 0.2868 0.2384 0.2364 0.2353
[…]
As we can see in Figures 1, 2, and Table 1, a 0.234 acceptance rate is still quite efficient in dimensions as low as 5. For an i.i.d. gamma target distribution, there is a significant decrease in the optimal acceptance rate when the dimension $d$ increases from 5 to 10. Subsequently, the optimal acceptance rate approaches 0.234 as the dimension increases. In the case of an i.i.d. beta target distribution, the acceptance rate is already close to 0.234 at $d=10$, and as the dimension increases further, the optimal acceptance rate observed in experiments gets closer to 0.234.
Our simulation study provides further evidence that for i.i.d. product target distributions, the optimal acceptance rate for maximizing ESJD is approximately 0.234 even in low dimensions, and that the optimal acceptance rate gets closer to 0.234 as the dimension increases. This result holds for both i.i.d. gamma and beta target distributions, as evidenced by the trace plots and histograms demonstrating the efficacy of the Metropolis algorithm across different dimensions.

### v3

@@ Abstract
For random-walk Metropolis (RWM) and parallel tempering (PT) algorithms, an asymptotic acceptance rate of around 0.234 is known to be optimal in certain high-dimensional limits. However, its practical relevance is uncertain due to restrictive derivation conditions. We synthesise previous theoretical advances in extending the 0.234 acceptance rate to more general settings, and demonstrate its applicability with a comprehensive empirical simulation study on examples examining how acceptance rates affect Expected Squared Jumping Distance (ESJD). Our experiments show the optimality of the 0.234 acceptance rate for RWM is surprisingly robust even in lower dimensions across various proposal, multimodal distributions that may not have an i.i.d. product density, and curved Rosenbrock target distributions with nonlinear correlation structure. Parallel tempering experiments also show that the idealized 0.234 spacing of inverse temperatures may be approximately optimal for low dimensions and non i.i.d. product target densities, and that constructing an inverse temperature ladder with spacings given by a swap acceptance of 0.234 is a viable strategy.
@@ 1 Introduction
Markov chain Monte Carlo (MCMC) methods, such as the random-walk Metropolis algorithm, are used to draw samples from complex high-dimensional target probability distributions. They enjoy strong theoretical asymptotic guarantees of accuracy, converging to the target distribution in stationarity [2, 28]. However, they may also be inefficient and prohibitively slow to provide a good approximation of the target distribution they are sampling [32]. Therefore, the analysis of their running time is an important practical issue to consider.
@@
[29] proved an important result for independently and identically distributed (i.i.d.) product target densities: in the state space high-dimensional limit $d\to\infty$, the asymptotic acceptance rate of a Metropolis algorithm is central to its efficiency. Under their assumptions, an asymptotic acceptance rate of approximately 0.234 maximises the efficiency of a random-walk Metropolis (RWM) algorithm with a Gaussian increment proposal distribution starting at stationarity. However, their assumptions on the target and proposals are very restrictive. Many practical scenarios often involve finite and relatively small dimensions. Additionally, many probability distributions do not have densities of the i.i.d. product form, and for those that do, sampling a one-dimensional target is much simpler than using RWM. A burgeoning active research area is dedicated to showing that this figure still applies beyond these assumptions, and much of this focuses on the i.i.d. product assumption.
@@
Many papers have shown that the optimal average acceptance rate of 0.234 for RWM holds greater generality than the assumptions used in Roberts et al. [29]. For more general theoretical results, Sherlock [38] replaces the restrictive i.i.d. product target and Gaussian proposal setting of Roberts et al. [29] with a set of “shell”, relative variability, and eccentricity conditions on the target, eccentricity condition on the jump distance matrix, and shell condition on the proposal. This framework does not require a product form and is based on the geometric properties of the (log) target distribution in high dimensions. Yang et al. [42] makes the 0.234 result applicable to targets with sparse dependency structures, such as those arising from Bayesian graphical models, using a set of sufficient conditions based on local dependencies and bounded asymptotic behaviour of derivatives. We expand on these results in Section 2. The 0.234 acceptance rate in optimal scaling has also been shown to apply to the spacing and swapping between inverse temperatures in the parallel tempering MCMC method [1, 36]. That being said, there are some cases in which 0.234 is not optimal, such as in the case of discontinuous targets [25] where the optimal acceptance rate is approximately 0.1353. Additionally, these results of optimality are still theoretical ideals, with required assumptions that could be further relaxed, and in the limit of $d\to\infty$ where the Markov chain converges to a diffusion process.
Other relaxations of the 0.234 assumptions include: non-Gaussian proposals [24], discrete hypercube target distributions [30], Gibbs random fields [3], perturbations from product targets using pre-conditioned proposals [5], independent target components with inhomogeneous scaling [6, 4, 33], partial updates where not all components are updated at once [23], elliptical symmetric unimodal target densities [39], and infinite-dimensional target distributions with non-trivial dependence structures [20].
@@
We begin our paper with a description of the RWM and its optimal scaling framework, as well as formally stating sufficient theoretical conditions for the 0.234 rule that are most relevant to our work in Section 2. In Section 3, we describe RWM experiments in low dimensions that show where an acceptance rate of 0.234 can be optimal in the sense that it maximises the Expected Squared Jumping Distance (ESJD) for various proposal and target densities which are not necessarily all i.i.d. product forms. Next, we describe the parallel tempering method and its own optimal scaling framework in Section 4, and show using experiments in Section 5 that the 0.234 swap acceptance rate in parallel tempering may also be optimal in lower dimensions for the multivariate Gaussian target density and a Gaussian mixture target density which is not of an i.i.d. product form. Lastly, we discuss the implications of our findings in Section 6.
@@ 2 Random-Walk Metropolis Background
We first introduce the Metropolis algorithm and the Optimal Scaling framework. We then discuss measures to evaluate the efficiency of MCMC algorithms, such as the expected squared jumping distance, and explain how the asymptotic acceptance rate of a MCMC algorithm is tied to the algorithm’s efficiency. Finally, we discuss some of the conditions for the 0.234 acceptance rate to be optimal that are most relevant to our work.
@@ 2.1 Metropolis Algorithm
A Metropolis algorithm [21] constructs a Markov chain $\mathbf{x}^{(0)},\mathbf{x}^{(1)},\dots$. If the chain is constructed to be irreducible and aperiodic, it is guaranteed to have the target distribution $\pi$ as its unique stationary distribution. This means that as the number of steps $t$ becomes large, the distribution of $\mathbf{x}^{(t)}$ converges to $\pi$. Therefore, after an initial burn-in period, the subsequent states of this single, long chain can be used as a sample from the target distribution $\pi$. When taking each step in the Markov chain, we generate a new state $\mathbf{y}=\mathbf{x}^{(t)}+\epsilon$ where $\epsilon$ is sampled from an increment “proposal” density $Q$ having $Q(\epsilon)=Q(-\epsilon)$ for all $\epsilon$ in its domain and accept this new state with probability $\alpha(\mathbf{x},\mathbf{y})$ given by $\alpha=\min\left\{1,\dfrac{\pi(\mathbf{y})}{\pi(\mathbf{x})}\right\}$, where, by a common abuse of notation, we let $\pi(\cdot)$ also denote the probability density function of the target distribution $\pi$. If the proposed value is accepted, we set the next state $\mathbf{x}^{(t+1)}=\mathbf{y}$. Otherwise, we set $\mathbf{x}^{(t+1)}=\mathbf{x}^{(t)}$.
@@ 2.2 Optimal Scaling Framework
Optimal Scaling [29, 33] is one of the most successful frameworks for performing asymptotic analysis of high-dimensional MCMC methods, and provides mathematically-grounded guidance on how to best optimise MCMC performance by tuning the “scaling” parameter(s) of the proposal distribution. Typically, a multivariate Gaussian distribution with mean 0 is used as the increment proposal distribution. The scaling of the Gaussian proposal is then given by a scaling factor $\sigma$, and this is often used as a variance $\sigma^{2}$ that is applied to the $d\times d$ identity matrix $I_{d}$. So, $\epsilon\sim\mathcal{N}(0,\sigma^{2}I_{d})$. If the proposal’s variance is too small, many proposed steps are accepted by the algorithm, but each step does not explore the state space much. Vice versa, if the proposal’s variance is too large, too few proposed steps are accepted, leading to slow exploration of the state space since the algorithm stays at a state for too long. Therefore, the proposal’s variance is crucial to the performance of the algorithm; the optimal scaling should balance the exploration of new areas in the state space with exploiting high-density areas of the target distribution. Different Metropolis algorithms have shown to have different optimal choices of the proposal distribution variance, $\sigma^{2}$.
@@
We next discuss how to evaluate the efficiency of an MCMC algorithm for determining optimal scaling. A very popular measure of efficiency both theoretically (e.g. Sherlock [38], Roberts and Rosenthal [36], Yang et al. [42]) and in practice (e.g. Pasarica and Gelman [26], Sherlock et al. [40]) is the Euclidean Expected Squared Jumping Distance (ESJD) metric which we use for our experiments. The expected squared jumping distance measures how far, in expectation, the MCMC chain moves in a single iteration. For the standard random-walk Metropolis algorithm, we define this as
$\mathbb{E}\left[\left\|\mathbf{x}^{(t+1)}-\mathbf{x}^{(t)}\right\|^{2}\right]\approx\frac{1}{n-1}\sum_{i=1}^{n-1}\left\|\mathbf{x}^{(t+1)}-\mathbf{x}^{(t)}\right\|^{2}$ (1)
@@
where $n$ is the total number of iterations of the algorithm. Maximising the ESJD aligns with minimizing the first-order auto-correlation of the Markov chain and subsequently maximises efficiency if the higher-order auto-correlations are monotonically increasing relative to the first-order auto-correlation [26, 42]. It is worth mentioning that ESJD primarily measures local convergence performance. Because heavy-tailed targets disrupt the monotonic relationship between first-order and higher-order autocorrelations, achieving global convergence for heavy-tailed targets requires maximising an adapted version of ESJD [17]. However, we do not consider heavy-tailed targets in this work, so the working definition of ESJD is sufficient.
@@
There are other notions of efficiency of a Markov chain, but in the high-dimensional limit $d\to\infty$, if the moment conditions in Assumption 2.1 holds, the chain converges to a diffusion process, and all efficiency measures are effectively equivalent [29, 33]. A key result of these referenced papers is that, given a Metropolis algorithm with a Gaussian increment proposal distribution $Q=\mathcal{N}(0,\frac{\ell^{2}}{d}I_{d})$ where $\ell>0$ is a fixed scaling constant and $I_{d}$ is the identity matrix, maximising the diffusion’s speed measure $h(\ell)$, which is a function of the scaling constant $\ell$, yields the most efficient asymptotic diffusion. Furthermore, the speed measure has a clear relation to a much simpler quantity to tune for: the asymptotic acceptance rate of the proposed new states (moves) of the algorithm, defined as
@@
Both the speed measure $h(\ell)$ and asymptotic acceptance rate $a(\ell)$ are functions of the scaling constant $\ell$. Given the restrictive conditions mentioned in Section 2.3, the scaling constant $\ell$ that maximises the speed measure $h(\ell)$ corresponds to an asymptotic acceptance rate $a(\ell)$ of approximately 0.234. Hence, an asymptotic acceptance rate of approximately 0.234 should necessarily optimise a measure of efficiency such as the ESJD of the algorithm.
2.3 Theoretical Assumptions for Optimality of a 0.234 Acceptance Rate
Certain work has shown that under restrictive assumptions, a 0.234 acceptance rate is asymptotically optimal for random-walk Metropolis algorithms in certain high-dimensional limits. We briefly list the theoretical assumptions of these works that are most relevant to our simulations. We first describe the seminal result of Roberts et al. [29] alluded to in the introduction in full.
Assumption Set 2.1.
(Roberts et al. [29]) Let $\mathbf{x}$ be a random variable from the target density $\pi$. The target distribution has an i.i.d. product density $\pi(\mathbf{x})=\prod_{i=1}^{d}f(x_{i})$. The component density $f:\mathbb{R}\to\mathbb{R}^{+}$ satisfies the following regularity conditions:
  1. 1.
    $f$ is positive and twice continuously differentiable on its support.
  2. 2.
    The derivative of its logarithm, $f^{\prime}/f$, is Lipschitz continuous.
  3. 3.
    The following moment conditions hold:
    $\mathbb{E}_{f}\left[\left(\frac{f^{\prime}(X)}{f(X)}\right)^{8}\right]<\infty\quad\text{and}\quad\mathbb{E}_{f}\left[\left(\frac{f^{\prime\prime}(X)}{f(X)}\right)^{4}\right]<\infty$ (2)
Under the conditions in Assumption 2.1, Roberts et al. [29] showed that for an RWM algorithm with a Gaussian proposal distribution starting in stationarity, the sped-up process converges to a limiting Langevin diffusion. Maximizing the speed of this limiting process, in turn, corresponds to tuning the algorithm to an asymptotic acceptance rate of approximately 0.234.
The Roberts et al. [29] result works because an i.i.d. product target looks locally the same no matter which coordinate you observe. Then, the analysis of Sherlock [38] shows that a similar homogeneity emerges for a much wider class of targets once the dimension is large.
Assumption Set 2.2.
(Sherlock [38]) Let $\mathbf{x}$ be a random variable from the target density $\pi$, $M(\mathbf{x})=||\nabla\log\pi(\mathbf{x})||$ be the norm of the gradient of the log target, and $H(\mathbf{x})=-\nabla^{2}\log\pi(\mathbf{x})$ be the negative Hessian matrix of the log target. The primary conditions are:
  1. 1.
    Target weak shell. There exist sequences $\{M^{(d)}\}$ and $\{H^{(d)}\}$, as $d\to\infty$:
[…]
  2. 2.
    Hessian relative variability. The Hessian is locally stable, meaning it does not vary significantly on the scale of a typical proposal jump. Formally, for any fixed $\mu>0$ and $\delta>0$, and for a standard Gaussian random vector $\mathbf{z}^{(d)}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$ independent of the target state $\mathbf{x}^{(d)}$, we require that as $d\to\infty$,
    $\mathbb{P}_{\mathbf{x,z}}\left(\frac{1}{H^{(d)}}\left|\left(\mathbf{z}^{(d)}\right)^{T}\Delta\left(\mathbf{x}^{(d)},t\mu\frac{M^{(d)}}{H^{(d)}}\mathbf{z}^{(d)}\right)\mathbf{z}^{(d)}\right|<\delta\quad\text{for all }t\in[0,1]\right)\to 1.$ (4)
  3. 3.
    Target eccentricity. The curvature is not disproportionately concentrated in any single direction. Denoting the eigenvalues of $\mathbf{H}(\mathbf{x})$ by $\beta_{i}(\mathbf{x})$:
[…]
  4. 4.
    Proposal shell. There exists a sequence $\{k_{u}^{(d)}\}$ such that the sequence of $d$-dimensional spherically symmetric proposals $\{Q^{(d)}\}$ satisfies this as $d\to\infty$:
    $\frac{||Q^{(d)}||}{k_{u}^{(d)}}\xrightarrow{m.s.}1.$ (6)
  5. 5.
    Jump distance matrix eccentricity. If the ESJD is defined with respect to a positive definite symmetric matrix $\mathbf{T}^{(d)}$ as $\mathbb{E}[(\mathbf{Y}^{(d)})^{T}\mathbf{T}^{(d)}\mathbf{Y}^{(d)}]$, then the eigenvalues of this jump distance matrix must be similarly well-behaved. Let $\tau_{i}(d)$ be the eigenvalues of $\mathbf{T}^{(d)}$. When $d\to\infty$:
    $\frac{\max_{i}\tau_{i}(d)}{\sum_{j=1}^{d}\tau_{j}(d)}\to 0.$ (7)
Put together, these geometric regularities recover the same weak-convergence proof as Roberts et al. [29] but with weaker assumptions on the target.
3 Random-Walk Metropolis Simulations
@@
In this section, we explore the applicability of the 0.234 acceptance rate figure for RWM algorithms using several practical RWM examples in lower dimensions. We begin with simple examples where the target densities meet all the key conditions in Roberts et al. [29] (provided as Assumption 2.1) required for the 0.234 acceptance rate to be optimal. Next, we investigate examples that do not use Gaussian proposal distributions, and then we explore other target densities that may lack these assumptions, such as smoothness or the i.i.d. product form. We estimate the Expected Squared Jumping Distance (ESJD) of the algorithm as a metric of the algorithm’s efficiency as mentioned in Section 2.2 and aim to see under what conditions an acceptance rate of approximately 0.234 maximises the ESJD for various target and proposal distributions in lower dimensions.
Our experimental procedure is as follows: we run RWM simulations using (except for Section 3.2) a Gaussian proposal distribution
$Q(\mathbf{y}\mid\mathbf{x})=\mathcal{N}(\mathbf{y}\mid\mathbf{x},\sigma^{2}I_{d})$ (8)
where $\sigma$ is a proposal scaling factor, or $\sigma^{2}$ can be thought of as a variance, and $I_{d}$ is the $d$-dimensional identity matrix. We run 40 different RWM simulations where each simulation has a different scaling factor, and each simulation has a burn-in of at least 1,000 steps before taking at least 200,000 steps. It is important to note that functions of acceptance rates are somewhat flat around the optimal values and that the values and figures reported are still prone to random error inherent in a Monte Carlo simulation. Therefore, we repeat this procedure over at least 20 computer seeds and average the results to reduce random error. This procedure is applied in each subsection and specific simulation details may be mentioned in a subsection if the experiment used more than these minimum values.
@@
We begin with simpler examples demonstrating that the 0.234 acceptance rate is still roughly optimal for lower dimensions with any i.i.d. product target distribution and a Gaussian proposal. As mentioned in Section 2, [29] proved the 0.234 optimal acceptance rate for i.i.d. product target distributions with a standard Gaussian proposal distribution in the infinite-dimensional limit, and [33] and [7] extended this with MCMC simulations showing that with a i.i.d. product target distribution and Gaussian proposal, the 0.234 acceptance rate seems approximately optimal for dimensions as low as 10. Our experiment provides further evidence for this being the case, and we extend this with simulations for $d=2,5$. We present two i.i.d. target densities $\pi_{1},\pi_{2}$ with different single-dimension component densities $f_{1},f_{2}$:
$\pi_{1}(\mathbf{x})$ $=\prod_{i=1}^{n}f_{1}(x_{i}),\quad f_{1}(x)=\text{Gamma}(x\mid 3,2),$ (9)
$\pi_{2}(\mathbf{x})$ $=\prod_{i=1}^{n}f_{2}(x_{i}),\quad f_{2}(x)=\text{Beta}(x\mid 3,2)$ (10)
The first example sets $f_{1}$ with shape parameter 3 and scale parameter 2, and the second example sets $f_{2}$ with shape parameters 3 and 2. We use a multivariate Gaussian with a mean of 0 as the increment proposal distribution in both cases. We run this experiment in dimensions $d\in\{2,5,10,30,50,100\}$.
Table 1: Empirical optimal RWM acceptance rates (i.e., maximizing ESJD) for i.i.d. product targets with a Gaussian proposal.
Target Density $d=2$ $d=5$ $d=10$ $d=30$ $d=50$ $d=100$
$\pi_{1}$: i.i.d. Gamma(3, 2) 0.3036 0.2378 0.2199 0.2101 0.2141 0.2140
$\pi_{2}$: i.i.d. Beta(3, 2) 0.3903 0.2937 0.2561 0.2319 0.2248 0.2159
The resulting ESJD-maximising acceptance rates are summarised in Table 1 and the i.i.d. Beta target has its trends of ESJD versus acceptance rate with different dimensions displayed in Figure 1. For the Gamma target, the optimal acceptance rate drops from 0.304 at $d=2$ to around 0.21 once $d\geq 10$; for the Beta target they fall from 0.390 to around 0.22 over the same range. Figure 1 confirms that the ESJD curves have a moving plateau towards a centre slightly below the classical 0.234 value as the dimension grows. Our simulation provides further evidence that the optimal acceptance rate for maximizing ESJD is approximately 0.234 even in low dimensions.
Figure 1: ESJD vs. acceptance rate for the i.i.d. Beta(3, 2) target distribution $\pi_{2}$ under RWM with a Gaussian proposal in dimensions $d\in\{2,5,10,30,50,100\}$ from top-left to bottom-right. Red dotted line indicates an acceptance rate of 0.234.
Since most ESJD versus acceptance rate plots look fairly similar and have a similar progression as dimension changes, the full set of curves for ESJD versus acceptance rate plots for other subsections are in Appendix A, including the results for the i.i.d. Gamma in Figure 9. As an aside, because the plateau widens appreciably in higher dimensions, the coarse 40-point grid pinpoints the optimum to approximately ±$\pm 0.01$.